OpenStudy (anonymous):

What is the exact area of a circle with a diameter of 37 inches

OpenStudy (maheshmeghwal9):

\[radius = \frac{diameter}{2}.\]\[Area=\pi r^2\] r=radius & \[\pi = 3.14 \space or \space \frac{22}{7}.\]

OpenStudy (anonymous):

area=pie*r^2 r = 37/2 = 18.5 inches Sub for r in the equation above.